OpenStudy (anonymous):

How do you solve this inequality? n-5/2<2n-6/3

OpenStudy (hunus):

is that \[\frac{n-5}{2}<\frac{2n-6}{3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Actually <_

OpenStudy (anonymous):

but I doubt it would make a difference

OpenStudy (hunus):

\[(\frac{n-5}{2})\le(\frac{2n-6}{3})\] \[(\frac{n-5}{2})(3)\le(\frac{2n-6}{3})(3)\] \[(\frac{n-5}{2})(2)(3)\le(2n-6)(2)\] \[(n-5)(3)\le(2n-6)(2)\] \[-15\le n-12\] \[ -3\le n\] \[n \ge -3\]

OpenStudy (anonymous):

THANK YOU SO MUCH! can you help me with one more?

OpenStudy (hunus):

Sure :)

OpenStudy (anonymous):

i'll post a question with your name and give another medal to you

OpenStudy (hunus):

okay :)